# LLM-Based Multi-Agent Systems for Software Engineering: Vision and the Road Ahead

### v1

@@ Abstract.
Integrating Large Language Models (LLMs) into autonomous agents marks a significant shift in the research landscape by offering cognitive abilities competitive to human planning and reasoning. This paper envisions the evolution of LLM-based Multi-Agent (LMA) systems in addressing complex and multi-faceted software engineering challenges. LMA systems introduce numerous benefits, including enhanced robustness through collaborative cross-examination, autonomous problem-solving, and scalable solutions to complex software projects. By examining the role of LMA systems in future software engineering practices, this vision paper highlights the potential applications and emerging challenges. We further point to specific opportunities for research and conclude with a research agenda with a set of research questions to guide future research directions.
@@
conference: International Workshop on Software Engineering in 2030; November 2024; Porto de Galinhas (Brazil)
@@ 1. Introduction
Autonomous agents, defined as intelligent entities that autonomously perform specific tasks through environmental perception, strategic self-planning, and action execution (Franklin and Graesser, 1996; Albrecht and Stone, 2018; Mele, 2001), have emerged as a rapidly expanding research field since the 1990s (Maes, 1993). Despite initial advancements, these early iterations often lack the sophistication of human intelligence (Unland, 2015). However, the recent advent of Large Language Models (LLMs) (Kasneci et al., 2023) has marked a turning point. This LLM breakthrough has demonstrated cognitive abilities nearing human levels in planning and reasoning (Achiam et al., 2023; Kasneci et al., 2023), which aligns with the expectations for autonomous agents. As a result, there is an increased research interest in integrating LLMs at the core of autonomous agents (Lo, 2023; Xi et al., 2023; Wang et al., 2023b) (for short, we refer to them as LLM-based agents in this paper).
Nevertheless, the application of singular LLM-based agents encounters limitations, since real-world problems often span multiple domains, requiring expertise from various fields. In response to this challenge, developing LLM-Based Multi-Agent (LMA) systems represents a pivotal evolution, aiming to boost performance via synergistic collaboration. An LMA system harnesses the strengths of multiple specialized agents, each with unique skills and responsibilities. These agents work in concert towards a common goal, engaging in collaborative activities like debate and discussion. These collaborative mechanisms have been proven to be instrumental in encouraging divergent thinking (Liang et al., 2023), enhancing factuality and reasoning (Du et al., 2023), and ensuring thorough validation (Wu et al., 2023b). As a result, LMA systems hold promise in addressing a wide range of complicated real-world scenarios across various sectors (Horton, 2023; Wang et al., 2023c, a), such as software engineering (Lo, 2023; Qian et al., 2023; Li et al., 2024; Hong et al., 2023).
@@
    Robustness and Fault Tolerance: LMA systems address robustness issues through cross-examination in decision-making, akin to code reviews and automated testing frameworks, thus detecting and correcting faults early in the development process. On their own, LLMs may produce unreliable outputs, known as hallucination (Zhang et al., 2023; Yang et al., 2023), which can lead to bugs or system failure in software development. However, by employing methods like debating, examining, or validating responses from multiple agents, LMA systems ensure convergence on a single, more accurate and robust solution. This enhances the system’s reliability, error handling, and aligns with best practices in software quality assurance.
@@ 2.2. LLM-based Autonomous Agent
Formally speaking, an LLM-based agent can be described by the tuple $\langle L,O,M,A,R\rangle$ (Cheng et al., 2024), where:
  • $L$ symbolizes the Large Language Model, serving as the agent’s cognitive core. It is usually equipped with extensive, potentially domain-specific, knowledge, enabling it to make informed decisions from observations, feedback, and rewards. Typically, an LLM possesses significant zero-shot or few-shot learning capabilities, and we interact with the LLM with prompts.
@@ 2.3. LLM-Based Multi-Agent Systems
In contrast to single-agent systems, a Multi-Agent system places a greater emphasis on coordination, communication, and knowledge sharing among agents. LMA systems are especially suited for complex tasks that require expertise across various knowledge domains. In this paper, we define that a Multi-Agent system, particularly an LLM-Based Multi-Agent system, includes two primary elements: an orchestration platform and LLM-based agents.
Orchestration Platform: The platform serves as the core infrastructure that manages interaction and information flow among agents. It is designed to ensure efficient operation and communication within the LMA system through several key functions: Agent Profiling: As we desire a broad spectrum of expertise, agents are defined with unique abilities and actions, fitting specific roles within the system. Profiles may be explicitly pre-defined (Qian et al., 2023; Hong et al., 2023) or dynamically generated by LLMs (Wang et al., 2023b).
Communication Mechanisms: Specifically, these communication mechanisms include:
  • Communication Paradigms: Styles of interaction, such as Cooperative (i.e., agents work together towards a shared goal) (Abdelnabi et al., 2023), Debate (i.e., agents engage in discussions to challenge and refine ideas) (Li et al., 2024), and Competitive (i.e., agents have conflicting objectives and compete against each other) (Wu et al., 2024).
  • Communication Structure: The organization of communication channels, which can be:
      Decentralized: Agents communicate directly with each other without a central controlling authority (Chen et al., 2023).
      Centralized: A central agent facilitates communication between all other agents (Agashe, 2023).
      Hierarchical: With clear layers of authority and responsibility, information flows from higher to lower levels (Zhao et al., 2024).
      Nested: Communication structures are embedded within one another, combining characteristics of hierarchical, decentralized, and centralized systems (Han et al., 2024).
  • Communication Content: The actual data exchanged among agents, often in text form. In the context of software engineering, this may be programming languages, commit messages (Zhou et al., 2023), Stack Overflow posts (He et al., 2022), etc.
LLM-based agents: We focus on the scenarios where agents within the Multi-Agent systems are powered with LLMs. Leveraging LLMs’ feedback integration capabilities, agents can engage in meaningful dialogues, enhancing cooperation and performance.
3. Current Development
In this section, we review the current applications of LMA agents in the software engineering domains.
Software Development. Several works have applied LMA systems in software development. Li et al. introduce a pioneering communicative Multi-Agent framework named role-playing (Li et al., 2024) with exceptional performance in code generation. ChatDev (Qian et al., 2023) divides the development process into designing, coding, testing, and documentation phases, with specialized teams of software agents, such as programmers and test engineers. Experiment results demonstrate that ChatDev can finalize the development of an application in less than seven minutes with a cost below one dollar. MetaGPT (Hong et al., 2023) embeds Standardized Operating Procedures (SOPs) (Belbin and Brown, 2022) into its workflow and further includes roles like product manager and QA Engineer. Subsequent studies have further explored the impact of role differentiation among LLM-based agents on collaborative software development  (Dong et al., 2023; Huang et al., 2023).
Frameworks. Numerous open-source repositories have been developed, which offer developers the means to customize and refine these agents for distinct tasks. AutoGen (Wu et al., 2023a) is an open-source framework that enables the creation of LMA applications through customizable, conversable agents that integrate LLMs, human inputs, and executable tools with flexible conversation patterns. Langroid (langroid contributors, 2024) implements its Multi-Agent pipeline based on the Actor Framework (Hewitt et al., 1973). Moreover, frameworks like LangChain (Chase, 2022), OpenAgents (Xie et al., 2023), AutoGPT (Significant Gravitas, [n. d.]) and GPT-Engineer (Osika, 2023), can also be adaptable for use in LMA systems, despite being originally designed on single-agent applications.

### v2

@@ Abstract.
Integrating Large Language Models (LLMs) into autonomous agents marks a significant shift in the research landscape by offering cognitive abilities competitive to human planning and reasoning. This paper explores the transformative potential of integrating Large Language Models into Multi-Agent (LMA) systems for addressing complex challenges in software engineering. LLMs enhance agents’ cognitive abilities, enabling more robust collaboration, autonomous problem-solving, and scalable project management. We review the current application of LMA systems in software engineering and discuss how LMA systems can improve software development through cross-examination, error reduction, and task automation. Additionally, the paper outlines the emerging challenges and presents a research agenda aimed at advancing LMA system capabilities in software engineering, offering key research questions to guide future exploration.
@@ 1. Introduction
Autonomous agents, defined as intelligent entities that autonomously perform specific tasks through environmental perception, strategic self-planning, and action execution (Franklin and Graesser, 1996; Albrecht and Stone, 2018; Mele, 2001), have emerged as a rapidly expanding research field since the 1990s (Maes, 1993). Despite initial advancements, these early iterations often lack the sophistication of human intelligence (Unland, 2015). However, the recent advent of Large Language Models (LLMs) (Kasneci et al., 2023) has marked a turning point. This LLM breakthrough has demonstrated cognitive abilities nearing human levels in planning and reasoning (Achiam et al., 2023; Kasneci et al., 2023), which aligns with the expectations for autonomous agents. As a result, there is an increased research interest in integrating LLMs at the core of autonomous agents (Lo, 2023; Xi et al., 2023; Wang et al., 2023c) (for short, we refer to them as LLM-based agents in this paper).
Nevertheless, the application of singular LLM-based agents encounters limitations, since real-world problems often span multiple domains, requiring expertise from various fields. In response to this challenge, developing LLM-Based Multi-Agent (LMA) systems represents a pivotal evolution, aiming to boost performance via synergistic collaboration. An LMA system harnesses the strengths of multiple specialized agents, each with unique skills and responsibilities. These agents work in concert towards a common goal, engaging in collaborative activities like debate and discussion. These collaborative mechanisms have been proven to be instrumental in encouraging divergent thinking (Liang et al., 2023), enhancing factuality and reasoning (Du et al., 2023), and ensuring thorough validation (Wu et al., 2023). As a result, LMA systems hold promise in addressing a wide range of complicated real-world scenarios across various sectors (Horton, 2023; Wang et al., 2023e, a), such as software engineering (Lo, 2023; Qian et al., 2024c; Li et al., 2024; Hong et al., 2023).
@@
    Robustness and Fault Tolerance: LMA systems address robustness issues through cross-examination in decision-making, akin to code reviews and automated testing frameworks, thus detecting and correcting faults early in the development process. On their own, LLMs may produce unreliable outputs, known as hallucination (Zhang et al., 2023; Yang et al., 2023b), which can lead to bugs or system failure in software development. However, by employing methods like debating, examining, or validating responses from multiple agents, LMA systems ensure convergence on a single, more accurate, and robust solution. This enhances the system’s reliability and aligns with best practices in software quality assurance.
@@ 2.2. LLM-based Autonomous Agent
Formally speaking, an LLM-based agent can be described by the tuple $\langle L,O,M,P,A,R\rangle$ (Cheng et al., 2024), where:
  • $L$ symbolizes the Large Language Model, serving as the agent’s cognitive core. It is equipped with extensive knowledge, potentially fine-tuned for specific domains, allowing it to make informed decisions based on observations, feedback, and rewards. Typically, an LLM suited for this role is trained on vast corpora of diverse textual data and comprises billions of parameters, such as models like ChatGPT [^1] https://openai.com/chatgpt/, Claude [^2] https://claude.ai/, and Gemini [^3] https://gemini.google.com/app. These models exhibit strong zero-shot and few-shot learning capabilities, meaning they can generalize well to new tasks with little to no additional training. Interaction with the LLM typically occurs through prompts, which guide its reasoning and responses.
@@
  • $P$ represents Perception, which represents the agent’s ability to sense, interpret, and understand its surroundings and inputs. This perception can involve processing structured and unstructured data from various sources such as text, visual inputs, or sensor data. Perception allows the agent to interpret the environment, transforming raw information into meaningful insights that guide decision-making and actions.
@@ 2.3. LLM-Based Multi-Agent Systems
A multi-agent system is a computational framework composed of multiple interacting intelligent agents that interact and collaborate to solve complex problems or achieve goals beyond the capability of any single agent (Wooldridge, 2009). These agents communicate, coordinate, and share knowledge, often bringing specialized expertise to address tasks across diverse domains.
With the integration of LLMs, LLM-Based Multi-Agent Systems have emerged. In this paper, we define that an LMA system comprises two primary components: an orchestration platform and LLM-based agents.
2.3.1. Orchestration Platform
The orchestration platform serves as the core infrastructure that manages interactions and information flow among agents. It facilitates coordination, communication, planning, and learning, ensuring efficient and coherent operation. The orchestration platform defines various key characteristics:
  1. (1)
    Coordination Models: Defines how agents interact, such as cooperative (collaborating towards shared goals) (Abdelnabi et al., 2023), competitive (pursuing individual goals that may conflict) (Wu et al., 2024), hierarchical (organized with leader-follower relationships) (Zhao et al., 2024), or mixed models.
  2. (2)
    Communication Mechanisms: Determines how the information flows between the agents: It defines the organization of communication channels, including centralized (a central agent facilitates communication (Agashe, 2023)), decentralized (agents communicate directly (Chen et al., 2023a)), or hierarchical (information flows through layers of authority (Zhao et al., 2024)). Moreover, it specifies the data exchanged among agents, often in text form. In software engineering contexts, this may include code snippets, commit messages (Zhou et al., 2023), forum posts (He et al., 2022, 2024a, 2024b), bug reports (Bettenburg et al., 2008), or vulnerability reports (Imtiaz et al., 2021).
  3. (3)
    Planning and Learning Styles: The orchestration platform specifies how planning and learning are conducted within the multi-agent system. It determines how tasks are allocated and coordinated among agents. It includes strategies like Centralized Planning, Decentralized Execution (CPDE) – planning is conducted centrally, but agents execute tasks independently, or Decentralized Planning, Decentralized Execution (DPDE) – both planning and execution are distributed among agents.
2.3.2. LLM-Based Agents
Each agent may have unique abilities and specialized roles, enhancing the system’s ability to handle diverse tasks effectively. Agents can be:
  1. (1)
    Predefined or Dynamically Generated: Agent profiles can be explicitly predefined (Hong et al., 2023) or dynamically generated by LLMs (Wang et al., 2023c), allowing for flexibility and adaptability.
  2. (2)
    Homogeneous or Heterogeneous: Agents may have identical functions (homogeneous) or diverse functions and expertise (heterogeneous).
Each LLM-based agent can be represented as a node $v_{i}$ in a graph $G(V,E)$, where edges $e_{i,j}\in E$ represent interactions between agents $v_{i}$ and $v_{j}$.
3. Literature Review
In this section, we provide a review of recent studies on LMA systems in software engineering. We categorize these applications according to the various stages of the software development lifecycle (SDLC), focusing on key areas such as requirements engineering, code generation, quality assurance, and software maintenance. Additionally, we highlight papers that explore the use of LMA systems for end-to-end software development, which encompasses multiple phases of the SDLC, rather than focusing on isolated stages.
3.1. Requirements Engineering
Requirements Engineering (Van Lamsweerde, 2000; Lo, 2024) focuses on defining and managing software system requirements. This discipline is divided into several key stages to ensure requirements meet quality standards and align with stakeholder needs. These stages include elicitation, modeling, specification, analysis, and validation (Hickey and Davis, 2004; Christel and Kang, 1992).
Elicitron (Ataei et al., 2024) is an LMA framework that focuses specifically on the elicitation stage. It utilizes LLM-based agents to represent a diverse array of simulated users. These agents engage in simulated product interactions, providing insights into user needs by articulating their actions, observations, and challenges. MARE (Jin et al., 2024) is another LMA framework that covers multiple phases of requirements engineering, including elicitation, modeling, verification, and specification. It employs five distinct agents performing nine different tasks to help generate high-quality requirements models and specifications.
3.2. Code Generation
Code generation is an essential task in software development, where source code is produced from high-level specifications, often written in natural language. In this subsection, we review the recent innovative LMA frameworks that enhance the process of code generation.
A prominent design approach in LMA systems involves role specialization and cross-validation (e.g., peer review) among agents (Li et al., 2023; Liu et al., 2023; Sun et al., 2024b; Olausson et al., 2023; Dong et al., [n. d.]). In this method, agents are assigned distinct roles, such as programmers or testers, and collaborate iteratively to enhance code quality. INTERVENOR (Wang et al., 2024b) pairs a Code Learner with a Code Teacher to iteratively improve code and fix bugs using compiler error feedback. Similarly, Self-repair (Olausson et al., 2023) involves a code agent generating programs and a feedback agent providing execution error explanations to iteratively improve code. AutoCoder (Lei et al., 2024) simulates coding tasks with a questioner and a programmer. Both agents cooperate and verify the execution through unit tests. SoA (Ishibashi and Nishimura, 2024) proposes to organize agents based on problem complexity. PairCoder (Zhang et al., 2024a) incorporates a Navigator agent for high-level planning and a Driver agent for specific implementation. The Navigator is responsible for proposing promising solution plans, selecting the current optimal plan, and directing the next iteration round based on execution feedback. The Driver follows the guidance of Navigator to undertake initial code generation, code testing, and refinement. MapCoder (Islam et al., 2024) replicates the problem-solving process in competitive programming through four specialized LLM agents. The Retrieval Agent retrieves examples of similar problems, guiding the solution process. The Planning Agent then creates a structured solution plan, assigning confidence scores to each approach. The Coding Agent follows the plan to generate and test code, while the Debugging Agent continues to correct errors by comparing the code with the initial plan.
CodeS (Zan et al., 2024) converts natural language requirements into fully functional code repositories using three specialized agents. The RepoSketcher defines the repository structure, the FileSketcher generates blueprints for each file, and the SketchFiller completes the code for each function. Cai et al. (Cai et al., 2023) present a framework to enhance human-LLM interaction through visual programming techniques. It allows users to design workflows for complex tasks using low-code interfaces, giving them more control over LLM outputs. The framework includes two agents: a Planning agent, which creates workflows, and an Executing agent, which follows the user-edited workflows. CodexGraph (Liu et al., 2024b) improves code generation, code understanding, and other software engineering tasks by navigating large code repositories. It builds graph databases using static analysis, extracting code symbols and their relationships. These symbols, represented as nodes, and their interactions, like inheritance and containment, form the edges of the graph. CodexGraph enables LLM-based agents to interact with these graph databases by generating and executing graph queries.
3.3. Software Quality Assurance
In this subsection, we review related work on testing, vulnerability detection, bug detection, and fault localization, with a focus on how LMA systems are being employed to enhance software quality assurance processes.
Testing. Fuzz4All (Xia et al., 2024) generates testing input for software systems across multiple programming languages. In this framework, a distillation agent reduces user input, while a generation agent creates and mutates inputs. AXNav (Taeb et al., 2024) is designed to automate accessibility testing. It interprets natural language test instructions and executes accessibility tests, such as VoiceOver, on iOS devices. AXNav includes a planner agent, an action agent, and an evaluation agent. WhiteFox (Yang et al., 2023a) is a fuzzing framework that tests compiler optimizations. It uses two LLM-based agents: one extracts requirements from source code, and the other generates test programs. Additionally, LMA systems are employed for tasks such as penetration testing (Deng et al., 2023), user acceptance testing (Wang et al., 2024d), and GUI testing (Yoon et al., 2024).
Vulnerability Detection. GPTLens (Hu et al., 2023) is an LMA framework for detecting vulnerabilities in smart contracts. The system includes LLM-based agents acting as auditors, each independently identifying vulnerabilities. A critic agent then reviews and ranks these vulnerabilities, filtering out false positives and prioritizing the most critical ones. MuCoLD (Mao et al., 2024) assigns roles like tester and developer to evaluate code. Through discussions and iterative assessments, the agents reach a consensus on vulnerability classification.
Bug Detection. Intelligent Code Analysis Agent (ICAA) (Fan et al., 2023) is used for bug detection in static code analysis. The agents have access to tools like web search, static analysis, and code retrieval tools. A Report Agent generates bug reports, while a False Positive Pruner Agent refines these reports to reduce false positives. Additionally, ICAA includes Code-Intention Consistency Checking, which ensures the code aligns with the developer’s intended functionality by analyzing code comments, documentation, and variable names.
Fault Localiztion. RCAgent (Wang et al., 2023b) performs root cause analysis in cloud environments by using LLM-based agents to collect system data, analyze logs, and diagnose issues. AgentFL (Qin et al., 2024) breaks down fault localization into three phases. The Comprehension Agent identifies potential fault areas, the Navigation Agent narrows down the codebase search, and the Confirmation Agent uses debugging tools to validate the faults.
3.4. Software Maintenance
In this subsection, we explore related work on debugging and code review, highlighting how LMA systems contribute to automating and improving software maintenance processes.
Debugging. Debugging involves identifying, locating, and resolving software bugs. Several frameworks, including MASAI (Arora et al., 2024), MarsCode (Liu et al., 2024a), AutoSD (Kang et al., 2023), and others (Tao et al., 2024; Ma et al., 2024; Chen et al., 2024), follow a structured process consisting of stages like bug reproduction, fault localization, patch generation, and validation. Specialized agents are typically responsible for each stage. FixAgent (Lee et al., 2024) includes a debugging agent and a program repair agent that work together to iteratively fix code by analyzing both errors and repairs. The system refines fault localization by incorporating repair feedback. The agents also articulate their thought process, improving context-aware debugging. The MASTER framework (Yang et al., 2024) employs three specialized agents. The Code Quizzer generates quiz-like questions from buggy code, the Learner proposes solutions, and the Teacher reviews and refines the Learner’s responses. AutoCodeOver (Zhang et al., 2024c) uses an agent for fault localization via spectrum-based methods, collaborating with others to refine patches using program representations like abstract syntax trees. SpecRover (Ruan et al., 2024) extends AutoCodeOver by improving program fixes through iterative searches and specification analysis, based on inferred code intent. ACFIX (Zhang et al., 2024b) targets access control vulnerabilities in smart contracts, focusing on Role-Based Access Control. It mines common RBAC patterns from over 344,000 contracts to guide agents in generating patches. DEI (Zhang et al., 2024f) resolves GitHub issues by using a meta-policy to select the best solution, integrating and re-ranking patches generated by different agents for improved issue resolution. RepoUnderstander (Ma et al., 2024) constructs a knowledge graph for a full software repository and uses Monte Carlo tree search to assist in understanding complex dependencies.
Code Review. Rasheed et al. (Rasheed et al., 2024a) developed an automated code review system that identifies bugs, detects code smells, and provides optimization suggestions to improve code quality and support developer education. This system uses four specialized agents focused on code review, bug detection, code smells, and optimization. Similarly, CodeAgent (Tang et al., 2024) performs code reviews with sub-tasks such as vulnerability detection, consistency checking, and format verification. A supervisory agent, QA-Checker, ensures the relevance and coherence of interactions between agents during the review process.
3.5. End-to-end Software Development
End-to-end software development involves managing the entire software development lifecycle, from gathering initial requirements to delivering the final product. In the context of LMA systems, each phase typically is handled by specialized agents which perform tasks suited to their expertise. The ultimate goal is to create fully autonomous systems capable of managing the entire development cycle. These systems often draw inspiration from established software development practices, such as Agile (Cohen et al., 2004), Scrum (Schwaber, 1997), and Waterfall (Petersen et al., 2009).
Pioneering works, such as ChatDev (Qian et al., 2024c) and MetaGPT (Hong et al., 2023), adopt a Waterfall model to automate the software development. ChatDev assigns agents as programmers, reviewers, and testers to handle design, coding, testing, and documentation. MetaGPT enhances the workflow by incorporating Standardized Operating Procedures (SOPs) (Belbin and Brown, 2022), with agents taking on roles like product managers and QA engineers. Recent advancements (Zhang et al., 2024d; Dong et al., 2023; Du et al., 2024) have further built on these designs, pushing the boundaries of fully autonomous development. Self-collaboration (Dong et al., 2023) presents a framework for improving code generation using multiple instances of ChatGPT. The approach simulates teamwork by assigning specific roles (analyst, coder, and tester) to distinct instances of ChatGPT. CTC (Du et al., 2024) involves multiple teams of LLM-based agents working together to generate high-quality software. Each team independently proposes solutions, and cross-communication allows them to refine their outputs by eliminating suboptimal proposals.
AgileCoder (Nguyen et al., 2024) and AgileGen (Zhang et al., 2024e) apply Agile methodologies in software development. AgileCoder (Nguyen et al., 2024) mirrors Agile’s role-based collaboration, assigning tasks like Product Manager and Developer to work in sprints. It also features a Dynamic Code Graph Generator, which updates code dependencies in real time, supporting context-aware generation and refinement. AgileGen (Zhang et al., 2024e) combines human-AI collaboration with Agile principles, using Gherkin for user acceptance criteria, ensuring alignment between requirements and code, and reducing user effort.
TGen (Mathews and Nagappan, 2024) explores Test-Driven Development (TDD), where test cases guide GPT-4 to generate correct code. LCG (Lin et al., 2024a) emulates various software process models, including Waterfall, TDD, and Scrum.
Additionally, several works (Rasheed et al., 2024b; Lin et al., 2024b; Qian et al., 2024b; Sami et al., 2024) adopt iterative and adaptive methodologies without explicitly detailing their models. For instance, ToP (Lin et al., 2024b) utilizes process models to guide agents in software development, which are refined through heuristic algorithms and process mining (Van Der Aalst, 2012). Qian et al. (Qian et al., 2024b) propose an iterative experience refinement framework that enables agents to continuously adapt to software development by leveraging prior experiences. This framework supports ongoing learning through the acquisition, utilization, and selective elimination of experiences. Co-Learning (Qian et al., 2024a) fosters collaboration through co-tracking, co-memorizing, and co-reasoning modules among agents. GameGPT (Chen et al., 2023b) automates and streamlines game development, employing various roles like game designers and code reviewers to enhance decision-making in game creation.